OpenStudy (anonymous):

write an equation 10 is multiplied by the sum of 3 and 2 . the product is divided by 2 and then 5 is added

OpenStudy (anonymous):

10 * 3 + 2 /2 + 5 just put the parenthesis is. (10 * (3+2))/2 + 5 that's how i would do it.

OpenStudy (anonymous):

thx

OpenStudy (mertsj):

\[\frac{10(3+2)}{2}+5\]

OpenStudy (anonymous):

evaluate 4-6(3-5)/3

OpenStudy (mertsj):

\[\frac{4-6(3-5)}{3} or 4-\frac{6(3-5)}{3}\]

OpenStudy (mertsj):

Which one is it?

OpenStudy (anonymous):

0 or 8

OpenStudy (mertsj):

Which problem is it? I posted two different possibilities for the problem as you posted it. Which one is it?

OpenStudy (anonymous):

the right one

OpenStudy (mertsj):

8

OpenStudy (anonymous):

thts wat i thought

OpenStudy (mertsj):

Good for you.

OpenStudy (anonymous):

6+5(4-3)^2

OpenStudy (anonymous):

11 right

OpenStudy (mertsj):

yes